OpenStudy (anonymous):

For the following system, use the second equation to make a substitution for y in the first equation. 3x + y = 2 y + 4 = 3x

OpenStudy (jackellyn):

solve for y in the second equation y = 3x -4 Put that in the first equation 3x + (3x-4) = 2 Simplify 3x + 3x - 4 = 2 6x -4 =2 Can you solve for x?

OpenStudy (anonymous):

x=1

OpenStudy (jackellyn):

So now you substitute that into any of the equations to find y. 3(1) + y=2 Can you solve for y?

OpenStudy (anonymous):

Oh they were just asking for 3x + 3x - 4 = 2

OpenStudy (jackellyn):

Well for extra practice and a medal can you tell me what y is?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y=-1? nt so sure

OpenStudy (jackellyn):

That is right @Hpbby! Now to check,you substitute both x and y into either equation. (-1) + 4 = 3(1) -1 + 4 = 3 3=3 3(1) + (-1)=2 3 -1 =2 2=2 Hooray! :)

OpenStudy (anonymous):

Thanks!! :)